# SPIN-ODE: Stiff Physics-Informed Neural ODE for Chemical Reaction Rate Estimation

### v2

@@ Abstract
Estimating rate coefficients from complex chemical reactions is essential for advancing detailed chemistry. However, the stiffness inherent in real-world atmospheric chemistry systems poses severe challenges, leading to training instability and poor convergence that hinder effective rate coefficient estimation using learning-based approaches. To address this, we propose a Stiff Physics-Informed Neural ODE framework (SPIN-ODE) for chemical reaction modelling. Our method introduces a three-stage optimisation process: first, a latent neural ODE learns the continuous and differentiable trajectory between chemical concentrations and their time derivatives; second, an explicit Chemical Reaction Neural Network (CRNN) extracts the underlying rate coefficients based on the learned dynamics; and third, fine-tune CRNN using a neural ODE solver to further improve rate coefficient estimation. Extensive experiments on both synthetic and newly proposed real-world datasets validate the effectiveness and robustness of our approach. As the first work on stiff Neural ODEs for chemical rate coefficient discovery, our study opens promising directions for integrating neural networks with detailed chemistry.
I Introduction
Neural Ordinary Differential Equations (Neural ODEs) have emerged as a powerful approach to bridging mathematical dynamical systems with neural networks. Their key advantage lies in representing system dynamics through continuous-depth neural networks. Neural ODEs have shown significant promise in the domains of physics and chemistry [1, 2, 3, 4]. However, despite these advances, their application to detailed chemical modelling remains limited in the computer science community, particularly in tasks such as reaction rate coefficient estimation and the discovery of novel chemical pathways [5].
@@ I Introduction
Understanding and predicting chemical reaction rate coefficients is a cornerstone of detailed chemical modelling. Unlike implicit modelling approaches [6, 7, 8, 9, 3] that focus on approximating system-level outputs, explicitly learning reaction rate coefficients demands a nuanced understanding of chemical kinetics. This task is particularly difficult in computer science, where knowledge transfer from domain-specific numerical chemistry to neural networks is still in its infancy. The fundamental obstacle lies in identifying and encoding the essential factors, such as temperature dependencies and chemical reaction paradigms, into a learnable and generalisable model architecture.
@@
In this work, we propose a novel framework that utilises the strengths of Neural ODEs to estimate reaction rate coefficients. As illustrated in Figure 1, our keen interest is to discover the correlations between chemical concentration trajectories and rate coefficients, so that the optimal rate coefficients can be found for ODE fitting. Rather than using neural networks as an implicit ODE system, our goal is to discover the hidden chemical parameters for detailed chemistry. We envision this integration of AI and detailed chemistry for scientific discovery, potentially uncovering new patterns, simplifying mechanisms, and offering interpretable insights into complex reaction systems. To summarise:
@@
  • We propose an explicit stiff physics-informed neural ODE framework (SPIN-ODE) for detailed chemistry. This framework efficiently models ODE trajectories while simultaneously extracting interpretable reaction rate coefficients.
@@
  • We introduce a three-stage optimisation strategy specifically designed to handle stiffness and enhance training stability. Our novel physical loss functions and trajectory resampling techniques significantly improve ODE fitting and rate coefficient estimation.
@@
  • Compared to existing approaches that primarily calibrate rate coefficients of stiff reactions based on initial close guesses, our method is the first work to automatically find optimal rate coefficients in the absence of prior knowledge.
II Related Work
@@
Our work is on detailed chemical kinetics via neural networks, which is closely related to neural networks and ODEs, and their applications in chemical modelling. Let us revisit those topics in detail.
II-A Numerical and Neural ODEs
@@
Numerical methods for solving ordinary differential equations (ODEs) have been extensively studied, with classical approaches such as Euler’s method, Runge-Kutta schemes, and multistep methods [12] forming the foundation of the field. These techniques aim to approximate solutions to initial value problems with varying degrees of accuracy, stability, and computational efficiency. Over the decades, significant advances have been made in adaptive step-size control, stiff equation solvers, and error estimation strategies. Recent developments have also explored the intersection of numerical analysis and machine learning, where neural networks and physics-informed approaches offer data-driven alternatives for solving ODEs. Neural ODEs were first introduced in [13], where backpropagation can be used to train an ODE model. Later on, many works [14, 15, 16, 17, 18, 19] followed up on this direction to further analyse and improve the performance. For example, [15] analyses the theory behind the infinite-depth neural ODEs and proposes adaptive-depth and data-controlled neural ODE paradigms. [18] demonstrates that a simple augmentation of the latent ODE space can achieve lower losses with few parameters and is more stable to train. To overcome the slow training process, [14] proposes a regularised neural ODE (RNODE) to learn well-behaved dynamics.  [20] proposes Fourier analysis to ODEs for temporal estimation and potentially high-order spatial gradients calculation. Later on, Neural ODEs are also applied to domain-specific applications, like time-series processing [16], latent space image processing [17], glucose dynamics [21] and robotics control [22].
@@
Meanwhile, a significant challenge to the current Neural ODEs is the stiff ODE system. [23] was the first to observe that explicit methods fail for certain chemical kinetic systems. The ”stiff“ chemical reactions can happen so fast that small errors can be magnified if the equations are integrated. Kim et al., [24] study on two classical stiff chemical systems using their proposed scaling and adjoint solution to mitigate stiffness. Fronk et al. present two consecutive studies [25, 26] to demonstrate both explicit and implicit solution to neural stiff ODE training. Physics-informed neural network (PINN) [27] is another solution by integrating more physical constraints to tackle the high-stiff regime. Ji et al. [28] apply steady state analysis in the physics-informed loss to reduce the stiffness of the reaction system. Seiler et al. [29] propose a method based on transfer learning and design a multi-head PINN network to divide and conquer the ODE system into low-stiff and high-stiff regions. Sulzer et al. [30] propose an autoencoder to solve stiff ODE in the latent space. The first- and second-order gradients are used as physical losses for optimisation.
II-B Chemical Neural Networks
@@
The rise of deep learning in the early 2010s has significantly expanded the scope of scientific discovery processes. Neural networks for chemical modelling have recently attracted significant research interest, like weather forecasting [1, 31, 32], air quality estimation [8, 6, 33, 7], and others. Chemical kinetics can be considered as a complex ODE system and researchers have been studying both implicit and explicit approaches for efficient chemical modelling. For example, Owoyele and Pal [3] propose to utilise neural ODE to simulate complete chemical kinetics implicitly for fast computation. DeepONet [2] integrates the recently proposed Neural Operator [34] into the neural network to exploit the dynamical structure of chemical reactions. Liu et al. [4] further improve it by proposing ChemODE, an attention based neural operator for fast and accurate chemical modelling, which can successfully estimate over hundreds of chemical reactions. Kolmogorov-Arnold Networks (KAN) were proposed recently [35] as an alternative to multi-layer perceptrons for general neural network applications. It attracts a lot of attention for chemical modelling [36]. The advantage of using KAN is that it provides fast neural scaling and ”white-box“ interpretability and generalisability. ChemKAN [9] was later proposed to learn hydrogen combustion chemistry modelling, achieving a twofold speed-up over detailed numerical chemistry.
@@
In contrast to implicit models that approximate the overall chemical system behaviour, explicit chemical neural networks aim to directly learn or represent the fine-grained chemical mechanisms themselves, such as reaction rates, activation energies, or reaction pathways. This explicit modelling is crucial for advancing scientific understanding, especially in domains where interpretability and physical consistency are as important as predictive performance. Despite its importance, relatively few studies have focused on using neural networks to capture detailed chemical kinetics in an interpretable and physically faithful manner. This gap is largely due to the complexity of chemical reaction networks and the challenge of integrating domain knowledge, such as thermodynamic constraints and reaction mechanisms, into machine learning models. Some notable efforts include ChemTab [5], which tabulates reaction rates using neural networks for combustion modelling, and ReactionNet [37], which predicts reaction rates by incorporating reaction templates and thermochemical data. Ji et al. [38] autonomously discover reaction pathways by regression on stoichiometric and rate coefficients. However, it is only used to calibrate the rate coefficients initialised with a small-scale uncertainty given fixed reaction pathways [39]. In this work, we argue that explicit neural representations of chemical reactions offer a promising direction for coupling accuracy, interpretability, and computational speed in scientific modelling, especially when integrated with recent developments in symbolic neural networks or physics-informed learning frameworks.
III Approach
@@
Figure 2: Our proposed SPIN-ODE framework for stiff reaction rate coefficient estimation. It includes three steps: 1) latent neural ODE fitting by optimising predicted ODE trajectories, 2) resampling data from step 1 and pre-training physics-based CRNN by optimising predicted time derivatives, and 3) fine-tuning CRNN with the ODE solver by optimising predicted concentration trajectories. The detailed structure of the CRNN is shown in the upper right corner.
III-A Preliminaries
@@
[…]
@@
in which $s_{i,j}^{f}$ and $s_{i,j}^{r}$ are the forward and reverse stoichiometric coefficients for species $j$ in reaction $i$, indicating molecules consumed and produced per reaction. The reaction rates are determined by a power-law expression involving the rate coefficient $k_{i}$ and reactant concentrations:
@@
Here we use []delimited-[][$\cdot$][ ⋅ ] to represent chemical concentrations. Reaction rate coefficient $k_{i}$ follows the Arrhenius relationship and is dependent on the temperature and molecular energy barriers, and is often determined experimentally. The overall rate of change for species concentration $Y_{j}$ is the sum of the consumption and production rates from all reactions:
@@
Therefore, the governing equations for the Robertson reaction system, shown in Eq (1), constitute a system of non-linear ODEs:
@@
Stiffness $\text{Stiffness ratio}=\frac{\max|Re(\lambda)|}{\min|Re(\lambda)|}.$ (6)
@@
2. Neural ODEs. Similar to chemical reactions, many systems across science and engineering domains can be represented by ODEs:
@@
where $y(t)$ is the system state variable at time $t$, and the function $f$ governs the system dynamic. Neural Ordinary Differential Equations (Neural ODEs) approximate unknown system dynamics $f$ using neural networks parametrised by $\theta$:
@@
Since observational data typically includes only discrete-time state variables $y(t)$, Neural ODEs integrate the approximated dynamics to predict system states at subsequent times:
[…]
@@
The loss function is defined as the distance, e.g., mean squared error (MSE), between model predictions and observed state variables. With the differentiable $ODESolver$, the gradient of the loss with respect to the network parameters $\theta$ can be calculated and leveraged to optimise the network using gradient-based algorithms. Chen et al. propose a widely-used approach for memory-efficient gradient calculation by solving the backward-in-time [13]. However, neural ODE training is extremely vulnerable to stiffness, where minor errors in the forward pass are rapidly amplified in the backward pass [24]. In this paper, we are specifically interested in solving stiff ODEs in atmospheric chemistry for rate coefficient estimation. Existing stiff neural ODE solutions, like collocation based ODE training [41, 42] and temporal regularization based ODE training [19] fail to fit ODE trajectories, and cannot extract the rate coefficients.
III-B Proposed method
We propose a novel neural network approach, SPIN-ODE, for stiff chemical modelling, which not only learns ODE fitting, but also explicitly extracts the rate coefficients for detailed chemistry. As shown in Figure 2, the whole network consists of three modules: MLPs, ODE solver, and Chemical Reaction Neural Network (CRNN). The overall process includes three steps: 1) latent ODE fitting, 2) resampling ODE trajectories for physics-based CRNN pre-training, and 3) fine-tuning CRNN with the ODE solver. The first step is to train MLPs to implicitly estimate ODE trajectories given input initial chemical concentrations. The second step is to resample from learned ODE trajectories to pre-train the CRNN for rate coefficient approximation. The final step is to refine rate coefficients by combining CRNN and an ODE solver for fine-tuning. Let us introduce them in detail.
SPIN-ODE step 1: Latent Neural ODE fitting In the first step, we fit the trajectory with a latent neural ODE network. We combine an MLP module and an ODE solver to implicitly learn the neural ODE curve fitting as Eq 9. Once it is trained, the MLP approximates the correlations between $y$ and $dy/dt$. The key component is the design of the denormalisation layer, which is crucial for ODE fitting in stiff chemistry. The motivation comes from [24], where the stiffness of the concentration roughly aligns with their time derivatives, such that they need to scale up to the same range for stable optimisation. Specifically, we describe the process as follows:

### v3

@@ Abstract
Estimating rate coefficients from complex chemical reactions is essential for advancing detailed chemistry. However, the stiffness inherent in real-world atmospheric chemistry systems poses severe challenges, leading to training instability and poor convergence, which hinder effective rate coefficient estimation using learning-based approaches. To address this, we propose a Stiff Physics-Informed Neural ODE framework (SPIN-ODE) for chemical reaction modelling. Our method introduces a three-stage optimisation process: first, a black-box neural ODE is trained to fit concentration trajectories; second, a Chemical Reaction Neural Network (CRNN) is pre-trained to learn the mapping between concentrations and their time derivatives; and third, the rate coefficients are fine-tuned by integrating with the pre-trained CRNN. Extensive experiments on both synthetic and newly proposed real-world datasets validate the effectiveness and robustness of our approach. As the first work addressing stiff neural ODE for chemical rate coefficient discovery, our study opens promising directions for integrating neural networks with detailed chemistry.
[…]
1405
1 Introduction
Neural Ordinary Differential Equations (Neural ODEs) have emerged as a powerful approach for representing system dynamics with neural networks. Their main advantage is the ability to model dynamics in continuous depth, offering a natural framework for time-dependent processes. They have shown significant promise in physics and chemistry [29, 35, 32]. However, despite these advances, their use in detailed chemical modelling remains limited in the computer science community, particularly for tasks such as reaction rate coefficient estimation and the discovery of new chemical pathways [41].”
@@ 1 Introduction
While implicit modelling approaches [1, 10, 17, 27, 35] approximate concentration trajectories, explicit estimation of rate coefficients [38] yields mechanistic insight and can be directly applied in atmospheric research, making our target fundamentally different. This task is particularly difficult in computer science, where knowledge transfer from domain-specific numerical chemistry to neural networks is still at an early stage. The fundamental obstacle lies in identifying and encoding the essential factors, such as temperature dependencies and chemical reaction paradigms, into a learnable and generalisable model architecture.
@@
In this work, we propose a novel framework that leverages the strengths of neural ODE to estimate reaction rate coefficients. As illustrated in Figure 1, our goal is to uncover the correlations between chemical concentration trajectories and rate coefficients, so that the optimal rate coefficients can be found by ODE fitting. Rather than using neural networks as an implicit ODE system, our goal is to discover the hidden rate coefficients for detailed chemistry. We envision this integration of AI and detailed chemistry for scientific discovery, potentially uncovering new patterns, simplifying mechanisms, and offering interpretable insights into complex reaction systems. To summarise:
@@
  • We propose an explicit stiff physics-informed neural ODE framework (SPIN-ODE) for detailed chemistry. This framework efficiently models ODE trajectories while simultaneously extracting interpretable reaction rate coefficients given reaction pathways.
@@
  • We introduce a three-stage optimisation strategy specifically designed to handle stiffness and enhance training stability. Our novel physical loss functions and trajectory resampling techniques significantly improve ODE fitting and rate constant estimation.
@@
  • Compared to existing approaches that primarily calibrate rate coefficients of stiff reactions based on close initial guesses, our method is the first work to automatically find optimal rate coefficients in the absence of prior knowledge.
2 Related Work
@@
Our study addresses detailed chemical kinetics with neural networks, building on advances in neural ODEs and their applications in chemical modelling.
2.1 Numerical and Neural ODEs
@@
Numerical methods for solving ordinary differential equations have been extensively studied, with classical approaches such as Euler’s method, Runge-Kutta schemes, and multistep methods [4] forming the foundation of the field. These techniques aim to approximate solutions to initial value problems with varying degrees of accuracy, stability, and computational efficiency. Over the decades, significant advances have been made in adaptive step-size control, stiff equation solvers, and error estimation strategies. Recent developments have also explored the intersection of numerical analysis and machine learning, where neural networks and physics-informed approaches offer data-driven alternatives for solving ODEs. Neural ODE was first introduced in [5], where backpropagation can be used to train an ODE model. Later on, many works [12, 34, 47, 46, 11, 15] followed up on this direction to further analyse and improve the performance. For example, [34] analyses the theory behind the infinite-depth neural ODEs and proposes adaptive-depth and data-controlled neural ODE paradigms. [11] demonstrates that a simple augmentation of the latent ODE space can achieve lower losses with few parameters and is more stable to train. To overcome the slow training process, [12] proposes a regularised neural ODE (RNODE) to learn well-behaved dynamics.  [30] proposes Fourier analysis to ODEs for temporal estimation and potentially high-order spatial gradients calculation. Later on, neural ODEs are also applied to domain-specific applications, like time-series processing [47], latent space image processing [46], glucose dynamics [49] and robotics control [21].
@@
Meanwhile, a significant challenge to the current neural ODEs is the stiff ODE system. [8] was the first to observe that explicit methods fail for certain chemical kinetic systems. Even when the observed concentration trajectories appear smooth, stiffness arises because some reactions evolve extremely fast while others are much slower, leading to hidden oscillatory dynamics and numerical instability. Kim et al. [25] propose stiff neural ODE, which introduces scaling and an adjoint solution strategy to mitigate stiffness, and demonstrate its effectiveness on two classical stiff chemical systems. Fronk et al. present two consecutive studies [14, 13] to demonstrate both explicit and implicit solution to stiff neural ODE training. Physics-informed neural network (PINN) [7] is another solution by integrating more physical constraints to tackle the high-stiff regime. Ji et al. [23] develop Stiff-PINN, which integrates Quasi-Steady-State-Assumptions (QSSA) into the network by predicting only non-QSS species, reconstructing QSS species algebraically, and constraining the physics-informed loss with the reduced system, thereby mitigating stiffness in chemical kinetics. Seiler et al. [42] present a method based on transfer learning and design a multi-head PINN network to divide and conquer the ODE system into low-stiff and high-stiff regions. Sulzer et al. [20] propose an autoencoder to solve stiff ODE in the latent space. The first- and second-order gradients are used as physical losses for optimisation.
2.2 Chemical Neural Networks
@@
The rise of deep learning in the early 2010s has significantly expanded the scope of scientific discovery processes. Neural networks for chemical modelling have recently attracted significant research interest, like weather forecasting [29, 2, 3], air quality estimation [17, 1, 48, 10], and others. Chemical kinetics can be considered as a complex ODE system and researchers have been studying both implicit and explicit approaches for efficient chemical modelling. For example, Owoyele and Pal [35] utilise neural ODE to simulate complete chemical kinetics implicitly for fast computation. Goswami et al. [16] leverages the recently proposed neural operator [33] to exploit the dynamical structure of chemical reactions. Liu et al. [32] further improve it by proposing ChemODE, an attention based neural operator for fast and accurate chemical modelling, which can successfully estimate over hundreds of chemical reactions. Kolmogorov-Arnold Networks (KAN) were presented recently [31] as an alternative to multi-layer perceptrons for general neural network applications. It has attracted significant attention for chemical modelling [26]. ChemKAN [27] was later introduced to learn hydrogen combustion chemistry modelling, achieving a twofold speed-up over detailed numerical chemistry.
@@
In contrast to implicit models that approximate the overall chemical system behaviour, explicit chemical neural networks aim to directly learn or represent the fine-grained chemical mechanisms themselves, such as activation energies, rate coefficients or reaction pathways. This explicit modelling is crucial for advancing scientific understanding, especially in domains where interpretability and physical consistency are as important as predictive performance. Despite its importance, relatively few studies have focused on using neural networks to capture detailed chemical kinetics in an interpretable and physically faithful manner. This gap is largely due to the complexity of chemical reaction networks and the challenge of integrating domain knowledge, such as thermodynamic constraints and reaction mechanisms, into machine learning models. Some notable efforts include ChemTab [41], which tabulates reaction rates using neural networks for combustion modelling, and ReactionNet [44], which predicts reaction rates by incorporating reaction templates and thermochemical data. Ji et al. [22] autonomously discover reaction pathways and rates by estimating stoichiometric matrix and rate coefficients. However, it is only used to calibrate the rate coefficients initialised with a small-scale uncertainty given fixed reaction pathways [43]. In this work, we argue that explicit neural representations of chemical reactions offer a promising direction for coupling accuracy, interpretability, and computational speed in scientific modelling, especially when integrated with recent developments in symbolic neural networks or physics-informed learning frameworks.
3 Approach
@@
Figure 2: Our proposed SPIN-ODE framework for stiff reaction rate coefficient estimation. It includes three steps: 1) black-box neural ODE fitting by optimising predicted ODE trajectories, 2) resampling data from step 1 and pre-training physics-based CRNN by optimising predicted time derivatives, and 3) fine-tuning CRNN by optimising predicted concentration trajectories. The detailed structure of the CRNN is shown in the upper right corner.
3.1 Preliminaries
@@
A $\rightarrow\textrm{B},$ (1)
$\textrm{B}+\textrm{B}$ $\rightarrow\textrm{C}+\textrm{B},$
$\textrm{B}+\textrm{C}$ $\rightarrow\textrm{A}+\textrm{C}.$
@@ A $\rightarrow\textrm{B},$ (1)
in which $s_{i,j}^{f}$ and $s_{i,j}^{r}$ are the forward and reverse stoichiometric coefficients for species $j$ in reaction $i$, indicating molecules consumed and produced per reaction. The reaction rates are determined by a power-law expression involving the rate coefficients $k_{i}$ and reactant concentrations:
@@
Here we use $\cdot$] to represent chemical concentrations. Reaction rate coefficients $k_{i}$ follows the Arrhenius relationship and is dependent on the temperature and molecular energy barriers, and is often determined experimentally. The overall rate of change for species concentration $Y_{j}$ is the sum of the consumption and production rates from all reactions:
@@
Therefore, the governing equations for the Robertson reaction system, shown in Eq. (1), constitute a system of non-linear ODEs:
@@
$\text{Stiffness ratio}=\frac{\max|Re(\lambda)|}{\min|Re(\lambda)|}.$ (6)
@@
2. Neural ODE. Similar to chemical reactions, many systems across science and engineering domains can be represented by ODEs:
@@
where $y(t)$ is the system state variable at time $t$, and the function $f$ governs the system dynamic. Neural ODE approximates unknown system dynamics $f$ using neural networks parametrised by $\theta$:
@@
Since observational data typically includes only discrete-time state variables $y(t)$, neural ODE integrates the approximated dynamics to predict system states at subsequent times:
$y(t_{1})$ $=y(t_{0})+\int_{t_{0}}^{t_{1}}NN_{\theta}(y(t),t)dt$ (9)
@@
The loss function is defined as the distance, e.g., mean squared error (MSE), between model predictions and observed state variables. With the differentiable $ODESolver$, the gradient of the loss with respect to the network parameters $\theta$ can be calculated and leveraged to optimise the network using gradient-based algorithms. Chen et al. propose a widely-used approach for memory-efficient gradient calculation by solving the backward-in-time [5]. However, neural ODE training is extremely vulnerable to stiffness, where minor errors in the forward pass are rapidly amplified in the backward pass [25]. In this paper, we are specifically interested in solving stiff ODEs in atmospheric chemistry for rate coefficient estimation. Existing stiff neural ODE solutions, like collocation based ODE training [19, 40] and temporal regularisation based ODE training [15] fail to fit ODE trajectories, and cannot extract the rate coefficients.
3.2 Proposed method
We propose a novel neural network framework, SPIN-ODE, for stiff chemical modelling, which not only learns ODE trajectories but also explicitly extracts reaction rate coefficients for detailed chemistry. As illustrated in Figure 2, the framework addresses stiffness through a three-step incremental learning strategy that combines the robustness of black-box stiff neural ODE with the physical interpretability of CRNN. First, a neural ODE is trained to fit concentration trajectories from initial chemical conditions. Second, the learned trajectories are resampled to pre-train the CRNN for rate coefficient estimation. Finally, the pre-trained CRNN is integrated with a stiff ODE solver to fine-tune the rate coefficients.
SPIN-ODE step 1: Trajectory fitting with a black-box neural ODE In the first step, we fit the trajectories using a black-box stiff neural ODE model, where the dynamics are learned by a simple multilayer perceptron (MLP) combined with an ODE solver. After training, the MLP approximates the mapping between $y$ and $dy/dt$. A key component is the design of the normalisation layers, which are essential for stable ODE fitting in stiff chemical systems. Our motivation follows  [25], which shows that stiffness arises from scale separations in concentrations and their time derivatives, and can be mitigated by scaling both to comparable ranges for stable optimisation. Specifically, we describe the process as follows: